OpenStudy (anonymous):

What is the fourth term in this multiplication pattern? multiplication pattern using the formula 2 • 6n – 1 A. 12 B. 72 C. 432 D. 2592

OpenStudy (anonymous):

are you asking what n is?

OpenStudy (anonymous):

if so what does this equal?

OpenStudy (anonymous):

shoot n is a exponent

OpenStudy (anonymous):

Im so stupid

OpenStudy (anonymous):

ok.. so please rewrite the Question. Im not sure what the question is. Are we solving for N? or what?

OpenStudy (anonymous):

looks like it is: \[2\times 6^n -1\] amd we meed to find the fourth term

OpenStudy (anonymous):

yay that

OpenStudy (anonymous):

Then sorry i cant help here/ I dont understand that. Sorreh @baconbaby1334

OpenStudy (anonymous):

or is it \[2\times 6^{(n-1)}\]???? beacause the answers dont match for the other form...

OpenStudy (anonymous):

yay I think it is that but I dont understand how that works

OpenStudy (anonymous):

for the fourth term we just need to replace n by 4 we get: \[2\times6^{(4-1)}=2\times6^{3}\] can you calculate that?

OpenStudy (anonymous):

432

OpenStudy (anonymous):

great! so it is answer C !

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw!